The total profit for a company in October was 16% higher than it was in September. The total profit for the two months was $507,600. Find the profit for each month.
Answer by nerdybill(5410)

The total profit for a company in October was 16% higher than it was in September. The total profit for the two months was $507,600. Find the profit for each month.
.
Let x = profit for September
then
1.16x = profit of October
.
x + 1.16x = 507600
x(1+1.16) = 507600
x(2.16) = 507600
x = 507600/2.16
x = $235,000 (September)
.
October:
1.16x = 1.16(235000) = $272,600 (October)